Find the number that is 7/12 less than its reciprocal
Answer by mananth(12270)

let the number be x
its reciprocal = 1/x
x= 1/x -7/12
x-1/x = -7/12
x^2- 1 /x = -7/12
12(x^2-1)= -7x
12x^2-12 = 7x
12x^2+7x-12 =0
use quadratic equation to find the roots
x1=(-7+sqrt(625))/24
x1= 0.75
x2=(-7-sqrt(625))/24
x2= -1.333
..
m.ananth@hotmail.ca